# thanks: Author contributions: J.E. and T. A. contributed equally to this workthanks: Author contributions: J.E. and T. A. contributed equally to this work

### v1

@@ Abstract
We focus on three dependent variables $(R_{d},V_{d},t_{d})$ accessible both experimentally and numerically, where the first ejected drop radius is $R_{d}$, $V_{d}$ is the velocity of the first ejected drop (at pinch off), and $t_{d}$ is the time between bubble rupture and jet pinch off as a function of the surfactant parameter $\beta$. In Fig. 2 the experimental points (solid) are colored by the bubble lifetime prior to bubble rupture. The shape of the numerical points (open) are given by $E$ and are colored by La. Fig. 2(a,b,c) show that as $\beta$ increases, the drop radius $R_{d}$ increases, drop velocity $V_{d}$ decreases, and pinch-off time $t_{d}$ increases for both experiments and simulations. Increasing $E$ or decreasing $La$ in the simulations results in $R_{d}$ and $t_{d}$ increasing and $V_{d}$ decreasing. The experiment shows a spread for each measured $\beta$. Such a spread is physical and arises from the variation in bubble lifetime, whose effect seems to be similar to increasing the excess surfactant $E$, which suggests that variations in accumulation of surfactant at the top of the bubble, as captured by the parameter $E$, may be an explanation for lifetime effects [51, 43, 52]. Indeed, when considering relationships implicit in $\beta$ (see Fig. 2(d,e,f)), the experimental and numerical points collapse onto a curve with strong agreement between the experimental and numerical data. Such agreement suggests that the jet dynamics are being resolved accurately by the simulation. Physically, the collapse onto a curve suggests that the jet dynamics may be independent of the surfactant; it is the cavity collapse that is modulated by the surfactant, as Marangoni stresses affect regions of high curvature where surfactant accumulates during the cavity collapse. Again, it should be noted that the exact degree of agreement between the simulation and experiment is sensitive to the particular model of the isotherm used.
@@
Figure 2: Effect of the surfactant parameter $\beta$ on the first ejected drop radius $R_{d}$, velocity $V_{d}$, and pinch-off time $t_{d}$. (a,b,c) Relationships between $(R_{d},V_{d},t_{d})$ and $\beta$, nondimensionalised by the bubble radius $R_{0}$, capillary velocity $\gamma_{c}/\mu$, and inertiocapillary timescale $\sqrt{\rho R_{0}^{3}/\gamma_{c}}$ respectively. (d,e,f) Relationships between non-dimensionalised $(R_{d},V_{d},t_{d})$. The same numerical and experimental data points are used throughout (a-f). For $\textit{Bo}=0.16$ as considered, $r_{\text{cap}}/R_{0}\approx 0.5$ and hence results with $E=0-1$ are shown.
@@
We show that the size amplification due to the surfactants occurs since Marangoni stresses disturb the self-similar collapse of the cavity shape, leading to a slower and thicker jet formation. Indeed, without surfactants, at low $\textit{La}=1000-5000$ [27], it has been shown that the interface profile is self-similar just before and after jet production [25, 26, 5, 27] at time $t_{0}$ with a $|t-t_{0}|^{\frac{2}{3}}$ scaling.
@@
Fig. 3 shows the cavity collapse ($t<t_{0}$) in physical (a) and self-similar (b) coordinates, and the jet production ($t>t_{0}$) in physical (c) and self-similar (d) coordinates. A typical clean bubble (gray) is compared to a corresponding contaminated bubble (colored) at low $\textit{La}=2400$. The clean case in Fig. 3 is the analogue to Figs. 3,4 of Lai et al. [27] ([27] did not consider surfactants).
@@
Figure 3: Self-similar profile comparison between a clean ($\textit{La}=2400,\textit{Bo}=0.16$) and correspondingly contaminated ($\textit{La}=2400,\textit{Bo}=0.16,\beta=0.3,\Delta\gamma_{\infty}=0.55,E=0.5$) bubble. The clean case is analogous to Figs. 3,4 of [27]. The interface is given by $z=h(r,t)$. (a) The time evolution of the cavity collapse in time $(t<t_{0})$ for the clean (gray) and contaminated bubble (colors). The times shown are $(t_{0}-t)/t_{ic}=0.018,0.024,0.03$, where $t_{ic}=\sqrt{\frac{\rho R_{0}^{3}}{\gamma_{c}}}$ is the inertia-capillary timescale. (b) The time evolution of the cavity collapse in self-similar coordinates, corresponding to curves in (a). (c) The time evolution of jet production ($t>t_{0}$) (same convention as (a)), at times $(t-t_{0})/t_{ic}=0.008,0.011,0.015,0.019$. (d) The time evolution of jet production in self-similar coordinates corresponding to curves in (c). In (b,d), $\min$ is taken over the profile of the cavity or jet. The colors for (a,b,c,d) are given by the surfactant concentration $\Gamma$ and Marangoni stress $d\gamma/ds$ for surface tension $\gamma$ and arclength $s$ measured from $r=0$. The particular choice of times plotted in (a,b,c,d) is the same as [27], which captures the self-similarity.
@@
As noted in the literature, for clean bubbles, at low La, the cavity forms a sharp corner prior to jet formation (see Fig. 3(a,b)), which results in high capillary pressure that leads to a fast and thin jet (see Fig. 3(c,d)). In comparison, as can be seen in Fig. 3(a,b), the case with surfactants has a smoother corner, which results in a smaller magnitude capillary pressure and leads to a slower and thicker jet (see Fig. 3(c,d)). Physically, there is an accumulation of surfactants (see Fig. 3(a,b)) near the corner of the cavity (such accumulation has also been seen in [10, 12] for high La), which results in a low surface tension region leading to Marangoni stress that smooths the corner. Mathematically, it has been shown that the possible self-similar interface profile [53] for inertio-capillary cavity collapse is a family parametrized by the far-field (with respect to the local region of cavity collapse) cone angle and interfacial flow field. Thus, the cavity shape can be interpreted as the result of Marangoni flow that acts to select a smoother corner.
@@
Separately, for $\textit{La}\gg 5000$, multiple capillary waves appear as precursor waves ahead of the dominant wave (referred to as the corner above). In such a regime, the details of the precursor waves play a key role in the jet behavior, and hence the drop amplification mechanism identified in this Letter, is not applicable. Indeed, previous studies have suggested that for $\textit{La}\gg 5000$, the Marangoni stress damps the capillary waves in such a way that the drop radius decreases [10, 11, 12, 13]; which we also observe in our numerical and experimental data (see Fig. 1(a) and supplementary material).
@@
As a direct extension, it would be interesting to include how other physical effects, such as surfactant solubility and viscoelasticity, would affect the trend reversal. Considering a larger class of non-uniform initial surfactant distributions may also be important. Finally, many Marangoni flow problems would likely benefit from the approach taken in this Letter, where experimentally-derived surface tension isotherms are used in the numerical simulations.
@@ Acknowledgements.
This work was supported by NSF grant 2242512 to L.D. This work used TAMU Faster at Texas A&M High Performance Research Computing through allocation OCE140023 to L.D. from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [54], which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.

### v2

@@ III Results
We further investigate the effect of surfactants on the first ejected jet drop for low La, systematically varying $\beta$ and comparing with the experimental case with 50% glycerol. Sensitivity to La and non-uniformity in the initial surfactant distribution is considered so as to compare with the experimental variations in viscosity and bubble lifetime. We focus on three dependent variables $(R_{d},V_{d},t_{d})$ accessible both experimentally and numerically, where the first ejected drop radius is $R_{d}$, $V_{d}$ is the velocity of the first ejected drop (at pinch off), and $t_{d}$ is the time between bubble rupture and jet pinch off as a function of the surfactant parameter $\beta$. In Fig. 3 the experimental points (solid) are colored by the bubble lifetime prior to bubble rupture. The shape of the numerical points (open) are given by $E$ and are colored by La. Fig. 3(a,b,c) show that as $\beta$ increases, the drop radius $R_{d}$ increases, drop velocity $V_{d}$ decreases, and pinch-off time $t_{d}$ increases for both experiments and simulations. Increasing $E$ or decreasing $La$ in the simulations results in $R_{d}$ and $t_{d}$ increasing and $V_{d}$ decreasing. The experiment shows a spread for each measured $\beta$. Such a spread is physical and arises from the variation in bubble lifetime, whose effect seems to be similar to increasing the excess surfactant $E$, which suggests that variations in accumulation of surfactant at the top of the bubble, as captured by the parameter $E$, may be an explanation for lifetime effects [poulain_ageing_2018, shaw_film_2024, auregan2025surface]. Indeed, when considering relationships implicit in $\beta$ (see Fig. 3(d,e,f)), the experimental and numerical points collapse onto a curve with strong agreement between the experimental and numerical data. Such agreement suggests that the jet drop production dynamics are being resolved accurately by the simulation. Again, it should be noted that the exact degree of agreement between the simulation and experiment is sensitive to the particular model of the isotherm used. The numerical collapse onto a single curve also suggests the following. Changing either $\beta$ or $E$ moves the points along the same curve in Fig. 3(d,e,f), even though the surfactant distribution on the jet changes. If the primary effect of the surfactants on jet drop production was through the effects of Marangoni stresses on the pinching of the jet after the jet production, then one would not expect such a collapse, due to the differing surfactant concentration. This suggests that the primary effect of the surfactants on jet drop production at low La is through the modification of the cavity shape (more precisely, the capillary pressure).
@@
Figure 3: Effect of the surfactant parameter $\beta$ on the first ejected drop radius $R_{d}$, velocity $V_{d}$, and pinch-off time $t_{d}$. (a,b,c) Relationships between $(R_{d},V_{d},t_{d})$ and $\beta$, nondimensionalized by the bubble radius $R_{0}$, capillary velocity $\gamma_{c}/\mu$, and inertiocapillary timescale $\sqrt{\rho R_{0}^{3}/\gamma_{c}}$ respectively. (d,e,f) Relationships between non-dimensionalized $(R_{d},V_{d},t_{d})$. The same numerical and experimental data points are used throughout (a-f). For $\textit{Bo}=0.16$ as considered, $r_{\text{cap}}/R_{0}\approx 0.5$ and hence results with E=01E=0-1 are shown.
@@
We show that the size amplification due to the surfactants occurs since Marangoni stresses disturb the self-similar collapse of the cavity shape, leading to a slower and thicker jet formation. Indeed, without surfactants, at low $\textit{La}=1000-5000$ [lai_bubble_2018], it has been shown that the interface profile is self-similar just before and after jet production [zeff_singularity_00, duchemin_jet_2002, ghabache_physics_2014, lai_bubble_2018] at time $t_{0}$ with a $|t-t_{0}|^{\frac{2}{3}}$ scaling.
@@
Fig. 4 shows the cavity collapse ($t<t_{0}$) in physical (a) and self-similar (b) coordinates, and the jet production ($t>t_{0}$) in physical (c) and self-similar (d) coordinates. A typical clean bubble (gray) is compared to a corresponding contaminated bubble (colored) at low $\textit{La}=2400$. The clean case in Fig. 4 is the analogue to Figs. 3,4 of Lai et al. [lai_bubble_2018] ([lai_bubble_2018] did not consider surfactants).
@@
Figure 4: Self-similar profile comparison between a clean ($\textit{La}=2400,\textit{Bo}=0.16$) and correspondingly contaminated ($\textit{La}=2400,\textit{Bo}=0.16,\beta=0.3,\Delta\gamma_{\infty}=0.55,E=0.5$) bubble. The clean case is analogous to Figs. 3,4 of [lai_bubble_2018]. The interface is given by $z=h(r,t)$. (a) The time evolution of the cavity collapse in time $(t<t_{0})$ for the clean (gray) and contaminated bubble (colors). The times shown are $(t_{0}-t)/t_{ic}=0.018,0.024,0.03$, where $t_{ic}=\sqrt{\rho R_{0}^{3}/\gamma_{c}}$ is the inertia-capillary timescale. (b) The time evolution of the cavity collapse in self-similar coordinates, corresponding to curves in (a). (c) The time evolution of jet production ($t>t_{0}$) (same convention as (a)), at times $(t-t_{0})/t_{ic}=0.008,0.011,0.015,0.019$. (d) The time evolution of jet production in self-similar coordinates corresponding to curves in (c). In (b,d), $\min$ is taken over the profile of the cavity or jet. The colors for (a,b,c,d) are given by the surfactant concentration $\Gamma$ and Marangoni stress $d\gamma/ds$ for surface tension $\gamma$ and arclength $s$ measured from $r=0$. The particular choice of times plotted in (a,b,c,d) is the same as [lai_bubble_2018], which captures the self-similarity.
@@
As noted in the literature, for clean bubbles, at low La, the cavity forms a sharp corner prior to jet formation (see Fig. 4(a,b)), which results in high capillary pressure that leads to a fast and thin jet (see Fig. 4(c,d)). In comparison, as can be seen in Fig. 4(a,b), the case with surfactants has a smoother corner, which in combination with a lower surface tension due to the presence of surfactants, results in a smaller magnitude capillary pressure and leads to a slower and thicker jet (see Fig. 4(c,d)). Physically, there is an accumulation of surfactants (see Fig. 4(a,b)) near the corner of the cavity (such accumulation has also been seen in [constante_dynamics_2021, pico_surfactantladen_2024] for high La), which results in a low surface tension region leading to Marangoni stress that smooths the corner. Mathematically, it has been shown that the possible self-similar interface profile [sierou_selfsimilar_2004] for inertio-capillary cavity collapse is a family parametrized by the far-field (with respect to the local region of cavity collapse) cone angle and interfacial flow field. Thus, the cavity shape can be interpreted as the result of Marangoni flow that acts to select a smoother corner. It could be valuable to consider the spatial and temporal evolution of the curvature in greater detail in the presence of surfactants (e.g., such as in [gordillo_capillary_2019]).
@@
Separately, for $\textit{La}\gg 5000$, multiple capillary waves appear as precursor waves ahead of the dominant wave (referred to as the corner above). In such a regime, the details of the precursor waves play a key role in the jet behavior, and hence the drop amplification mechanism identified in this paper, is not applicable. Indeed, previous studies have suggested that for $\textit{La}\gg 5000$, the Marangoni stress damps the capillary waves in such a way that the drop radius decreases [constante_dynamics_2021, pierre_influence_2022, pico_surfactantladen_2024, vega_influence_2024]; which we also observe in our numerical and experimental data (see Fig. 2(a) and supplementary material).
IV Conclusion
@@
As a direct extension, it would be interesting to include how other physical effects, such as surfactant solubility and viscoelasticity, would affect the trend reversal. Considering a larger class of non-uniform initial surfactant distributions, may also be important. Finally, there are many problems where the details of Marangoni flow are important, including surfactant spreading [gaver_dynamics_1990], coating flows [quere_fluid_1999], bubble and droplet breakup, surface and breaking waves [erinin_comparison_2023] Many Marangoni flow problems would likely benefit from the approach taken in this paper, where experimentally-derived surface tension isotherms are used in the numerical simulations.
Rodríguez-Aparicio et al. [rodriguez_critical_2025] have posted an independent study that reports experimental results at low La compatible with the discussion presented here.
The code for the main text is available online [eshima_code_2025]. The numerical and experimental data for the main text is available online [eshima_data_2025].
@@ Acknowledgements.
Contribution: Conceptualization: JE, TA, LD. Investigation: JE, TA, PKF, SP, HAS, LD; experiments were performed by TA, while numerical simulations were performed by JE. Formal Analysis: JE, TA, LD, PKF, HAS, SP. Software: JE, PKF, SP; the surfactant transport module was developed by PKF with the basilisk library developed by SP. Writing – Original Draft Preparation: JE, TA, LD. Writing – Review & Editing: JE, TA, LD, PKF, SP, HAS.
This work was supported by NSF grant 2242512 to L.D. This work used TAMU Faster at Texas A&M High Performance Research Computing through allocation OCE140023 to L.D. from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [boerner_access_2023], which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.